OpenStudy (anonymous):

What is the product in lowest terms? -5/12 x 8/13 PLEASE HELP QUICK

OpenStudy (usukidoll):

multiply the fractions out and simplify

OpenStudy (anonymous):

\[-5/12 * 8/13 = -40/156\] Both are divisible by 4, so divide numerator and denominator by 4 \[(-40/4)/(156/4) = -10/38\] Both are divisible by 2, so divide numerator and denominator by 2 \[(-10/2)/(38/2) = -5/19\]